OpenStudy (anonymous):

Boolean algebra: Factor to obtain a product of sums (answer should have 3 terms): WXY + WX'Y + WYZ + XYZ' Please walk me through it? Show me which rules you apply?

OpenStudy (helder_edwin):

\[ \large WXY+WX'Y+WYZ+XYZ'= \] \[ \large =WY(X+X')+WYZ+XYZ' \] \[ \large =WY1+WYZ+XYZ' \] \[ \large =WY(1+Z)+XYZ'=WY1+XYZ'=WY+XYZ' \]